OpenStudy (lexi724):

MEDAL/FAN: MATH HELP! A car is traveling at a rate of 62 kilometers per hour. What is the car's rate in kilometers per minute? How many kilometers will the car travel in 2 minutes?

OpenStudy (lexi724):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

this looks like a conversion problem

HanAkoSolo (jamierox4ev3r):

there are two ways that this can be done. So for now, I'll just give you the easier one

HanAkoSolo (jamierox4ev3r):

Since the unit of kilometers stays the same, you would set this up as a ratio problem

OpenStudy (lexi724):

Please do haha I have 3 months of math to finish by dec 18

HanAkoSolo (jamierox4ev3r):

wow, that sounds painful

OpenStudy (lexi724):

yeah while working 2 jobs a 9-5and a 5:30 -9:30 5 days a week it is

OpenStudy (lexi724):

and my only 2 days off are class days D:

HanAkoSolo (jamierox4ev3r):

Alright. So far, we know that there is a rate of 62k per hour. so we start off with this: \(\Large\frac{62 km}{hour}\)

OpenStudy (lexi724):

62/2=31

HanAkoSolo (jamierox4ev3r):

since we know that there are 60 minutes in 1 hour, we multiply this onto the existing fraction. \(\Large\frac{62~km}{1 ~hour}\times\frac{1 hour}{60 minutes}\)

OpenStudy (lexi724):

62*60*1

OpenStudy (lexi724):

3720

HanAkoSolo (jamierox4ev3r):

since the "hour" units are the same, we can cancel these out, which gives us: \(\Large\frac{62~k}{60~minutes}\)

HanAkoSolo (jamierox4ev3r):

if we simplify this down you get 31/30. I'm not sure what you were doing, but the values that you are obtaining are much too high

OpenStudy (lexi724):

i cross multiplied

HanAkoSolo (jamierox4ev3r):

recall that if the car is going at 62 kilometers per hour, the it will be going less per minute, given that the speed stays the same. we want to maintain equivalent values

OpenStudy (lexi724):

1.0333333

HanAkoSolo (jamierox4ev3r):

with these types of problems, and when adding the units, you don't need to cross multiply. and yes, 1.0333333 is what you should be getting

OpenStudy (lexi724):

Thank you!! i will run to you often feel free to deny help i mostly need help setting up then i can solve

HanAkoSolo (jamierox4ev3r):

no problem. And best of luck! It must be difficult to work with such a demanding schedule. Keep studying hard!